OpenStudy (anonymous):

For questions 3–4, solve the equation using square roots. x2 – 81 = 0 (1 point) mc003-2.jpg –81, 81 –9, 9 no real number solutions

OpenStudy (anonymous):

\[x^2-81=0\\ x^2=81\\ x=\pm\sqrt{81}\\ x=\pm9\]